OpenStudy (adovbush):

Simplify. 9 – (3)(–5) A. 24 B. 30 C. –30 D. –6

OpenStudy (cggurumanjunath):



OpenStudy (adovbush):

A?

OpenStudy (xaze):

9-(-15) = 9 + 15 =?

OpenStudy (adovbush):

24 Thank you.

OpenStudy (xaze):

^^